# Lambda Expected Shortfall

### v1

@@ 4 Dual representation
$\mathrm{ES}_{\Lambda}(X)$ $=\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right)$ [Theorem 2]
@@
$=\sup_{\mathbb{Q}\in\mathcal{M}_{1,f}}\sup_{x\in\mathbb{R}}\left\{\mathbb{E}_{\mathbb{Q}}[X]\wedge x~:~\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}},~\mathbb{P}\mbox{-almost surely}\right\},$ [definition of $\mathcal{P}_{\Lambda(x)}$]
@@
$\mathbb{Q}\left(\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\right)=1$ $\iff\mathbb{E}_{\mathbb{P}}\left[\mathds{1}_{\left\{\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}}\right\}}\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\right]=1$
@@
$=\sup_{y\in[a,b]}\inf_{\mathbb{E}_{\mathbb{Q}}[Y]=t}\left(\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right)$ [minimax theorem (Fan, 1953)]
$=\sup_{y\in[a,b]}\left(\inf_{\mathbb{E}_{\mathbb{Q}}[Y]=t}\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right)$ [exchangeability of two infimums]
@@ 5.1 Representing Lambda ES as a minimization
If $\phi$ is further cash additive (naturally defined for interval-valued functions), then we call $\rho$ a Bayes risk measure, and $\phi$ the corresponding Bayes estimator. It is clear that $(\mathrm{VaR},\mathrm{ES})$ is a Bayes pair by (3). A natural question is whether we can also write $\mathrm{ES}_{\Lambda}$ as the minimum of some loss function and find its corresponding minimizer. Ideally, we may expect to find the relation between $\mathrm{VaR}_{\Lambda}$ and $\mathrm{ES}_{\Lambda}$ similar to $(\mathrm{VaR},\mathrm{ES})$ in optimization. Below we show a representation of $\mathrm{ES}_{\Lambda}$ based on the relation (3), which we call the RU representation of $\mathrm{ES}_{\Lambda}$ for simplicity. Define the mapping $T_{\Lambda}:\mathbb{R}\times\mathbb{R}\times L^{\infty}\to(-\infty,\infty]$ as
@@
In the next result, we also study convexity of $T_{\Lambda}$ with respect to different variables. We use the term “joint convexity” when there is more than one variable to emphasize that the property is different from convexity in each variable.
@@
(ii) The “if” part is clear by the convexity of $x\mapsto 1/(1-\Lambda(x))$. To show the “only if” part, suppose (21) is convex in $x$. Let
$\bar{x}=\inf\{x\in\mathbb{R}:\Lambda(x)=0\}.$
Right-continuity of $\Lambda$ yields that $\Lambda(\bar{x})=0$. We first prove $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in(-\infty,\bar{x})$. Suppose for contradiction that
@@
Next, we show that $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in\mathbb{R}$. Because $x\mapsto 1/(1-\Lambda(x))$ is decreasing, it suffices to show that $x\mapsto 1/(1-\Lambda(x))$ is continuous at $\bar{x}$ if $\bar{x}<\infty$. Suppose for contradiction that $\Lambda(\bar{x}-)>0$. Because of (24), there exists $a_{1}\in(-\infty,\mathrm{ess\mbox{-}sup}(X))$, such that
$\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]>\bar{x}.$
It is clear that $\mathbb{E}[(X-a_{1})_{+}]>0$ and thus
@@
$=\frac{\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]}{2}+\frac{\mathbb{E}\left[a_{1}+(X-a_{1})_{+}\right]\vee\bar{x}}{2}$
$<\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]$
$=\lim_{x\uparrow\bar{x}}\left\{\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda((x+\bar{x})/2)}(X-a_{1})_{+}\right]\vee\frac{x+\bar{x}}{2}\right\}.$
@@
$\mathbb{E}\left[\frac{a+b}{2}+\frac{1}{1-\Lambda((x+y)/2)}\left(X-\frac{a+b}{2}\right)_{+}\right]\vee\frac{x+y}{2}$
$=\left\{\frac{a+t}{2}+\frac{1-\Lambda(x)}{1-\Lambda((x+y)/2)}\left(t-\frac{a+t}{2}\right)\right\}\vee\frac{x+y}{2}>t.$
@@
For the RU representation of $\Lambda$-ES, Theorem 4 indicates that we do not guarantee joint convexity (or quasi-convexity) of the objective (21) in $(a,x)\in\mathbb{R}^{2}$ or $(a,x,X)\in\mathbb{R}\times\mathbb{R}\times L^{\infty}$ unless $\Lambda$ is a constant (i.e., $\mathrm{ES}_{\Lambda}$ is an ES). Nevertheless, the mapping (21) is convex in $x\in\mathbb{R}$ when the function $x\mapsto 1/(1-\Lambda(x))$ is convex. We give two examples satisfying this condition: $\Lambda$ is a constant (corresponding to the usual ES) and $\Lambda(x)=(e^{ax}+1)^{-1}$ for a>0a>0.444The example $\Lambda(x)=(e^{ax}+1)^{-1}$, $a>0$, $x\in\mathbb{R}$, provides a suggestion of a non-trivial choice of the $\mathrm{ES}_{\Lambda}$ risk measure to use in practice. By Theorem 4, an obvious advantage of such a choice is that it makes the objective of a practical $\Lambda$-ES-based optimization problem convex in the variable $x$. Moreover, Theorem 4 shows that $\mathrm{ES}_{\Lambda}$ has a similar feature to a Bayes risk measure, as it can be represented as the minimum of an expected loss function with an additional transformation. The corresponding minimizer is the interval of the left- and right-quantiles at the level of $\Lambda(x^{*})$ instead of $\mathrm{VaR}_{\Lambda},\mathrm{VaR}^{+}_{\Lambda}$].
@@ 5.2 Optimization with Lambda ES
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. Based on the theoretical results in the previous sections, we demonstrate general ideas of solving optimization problems with $\Lambda$-ES as a constraint or an objective. For $n\in\mathbb{N}$, let $\mathbf{L}=(L_{1},\dots,L_{n})\in(L^{\infty})^{n}$ represent a vector of losses, $\bm{\theta}\in\Theta$ represent a decision variable, where $\Theta$ is a convex set of actions, and $f:\Theta\times\mathbb{R}^{n}\to\mathbb{R}$ represent a loss function. For a typical example in finance, we can use $\mathbf{L}$ as the vector of losses from multiple assets and $\bm{\theta}$ as a portfolio weight vector.
First, we are interested in a problem where the decision maker aims at minimizing an objective risk measure $\rho:L^{\infty}\to\overline{\mathbb{R}}$ of the aggregate loss $f(\bm{\theta},\mathbf{L})$, guaranteeing that the $\Lambda$-ES of the total loss does not exceed a pre-specified value $\ell\in\mathbb{R}$. Namely, we consider the following optimization problem:
@@
$\iff\mbox{for all }x\in\mathbb{R}:~\mathrm{ES}_{\Lambda(x)}(f(\bm{\theta},\mathbf{L}))\wedge x\leqslant\ell$
@@
Proposition 4 implies that an optimization problem with a $\Lambda$-ES constraint below level $\ell\in\mathbb{R}$ can be equivalently converted to a problem with the constraint on $\mathrm{ES}_{\Lambda(\ell)}$ below the same level $\ell$. ES-constrained optimization problem has been studed extensively in the literature (see e.g., Krokhmal et al., 2002).
Another natural question to study is how to minimize $\Lambda$-ES as an objective. We consider the following optimization problem:
@@
In the current work, we aim to provide general insights into solving $\Lambda$-ES-based optimization problems, as stated in the result below. In specific problems, we may also consider some constraints along with the problem (27), which does not change the nature of our equivalence result. Define the mapping $\widetilde{T}_{\Lambda}:\Theta\times\mathbb{R}\times\mathbb{R}\times(L^{\infty})^{n}\to\overline{\mathbb{R}}$ as
@@ 6.2 Dominance of ES and Lambda ES
One can check that $\rho$ is quasi-convex, law-invariant, and $\rho\geqslant\mathrm{VaR}^{+}_{0}$. However, the condition $\rho\geqslant\mathbb{E}$ fails. Therefore, $\mathbb{E}$ is not the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}^{+}_{0}$ and thus (8) in Theorem 1 fails for $\alpha=0$.
@@
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (30)
@@
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (31)
@@
$\mathrm{ES}_{\alpha}=\min\{\rho:\underline{L}^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (32)
@@
The proofs for (30), (31), and the case of $\alpha\in(0,1)$ for (32) follow directly from those for Theorem 1 whose arguments still hold on $L^{1}$. We only need to prove (32) for $\alpha=0$.
@@
The proof of the following result follows the same arguments as that for Theorem 2, extending the space from $L^{\infty}$ to $L^{1}$ or $\underline{L}^{1}$.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1]$, the smallest quasi-convex and law-invariant risk measure on $L^{1}$ dominating $\mathrm{VaR}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR_Λ and $\rho$ is quasi-convex and law-invariant}.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, the smallest quasi-convex and law-invariant risk measure on $L^{1}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law-invariant}.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to[0,1)$, the smallest quasi-convex and law-invariant risk measure on $\underline{L}^{1}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law-invariant}.
@@
6.3 Dual and RU representations for Lambda ES
In this section, we restate the dual representation (Theorem 3) and the RU representation (Theorem 4) for $\Lambda$-ES on $L^{1}$. The proofs of both results below follow from the same arguments as those for Theorems 3 and 4 by replacing $L^{\infty}$ with $L^{1}$.

### v2

@@ 4 Dual representation
$\mathrm{ES}_{\Lambda}(X)$ $=\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right)$
@@
$=\sup_{\mathbb{Q}\in\mathcal{M}_{1,f}}\sup_{x\in\mathbb{R}}\left\{\mathbb{E}_{\mathbb{Q}}[X]\wedge x~:~\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}},~\mathbb{P}\mbox{-almost surely}\right\},$
@@
$\mathbb{Q}\left(\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\right)=1$ $\iff\mathbb{E}_{\mathbb{P}}\left[\mathbf{1}_{\left\{\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}}\right\}}\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\right]=1$
@@
$=\sup_{y\in[a,b]}\left(\inf_{\mathbb{E}_{\mathbb{Q}}[Y]=t}\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right)$ [minimax theorem (Fan, 1953)]
@@ 5.1 Representing Lambda ES as a minimization
If $\phi$ is further cash additive (naturally defined for interval-valued functions), then we call $\rho$ a Bayes risk measure, and $\phi$ the corresponding Bayes estimator. It is clear that $(\mathrm{VaR},\mathrm{ES})$ is a Bayes pair by (5). A natural question is whether we can also write $\mathrm{ES}_{\Lambda}$ as the minimum of some loss function and find its corresponding minimizer. Ideally, we may expect to find the relation between $\mathrm{VaR}_{\Lambda}$ and $\mathrm{ES}_{\Lambda}$ similar to $(\mathrm{VaR},\mathrm{ES})$ in optimization. Below we show a representation of $\mathrm{ES}_{\Lambda}$ based on the relation (5), which we call the RU formula of $\mathrm{ES}_{\Lambda}$. Define the mapping $T_{\Lambda}:\mathbb{R}\times\mathbb{R}\times L^{\infty}\to(-\infty,\infty]$ as
@@
The next result also demonstrates the convexity of $T_{\Lambda}$ in different variables. We use the term “joint convexity” when there are multiple variables to emphasize its difference from marginal convexity.
@@
(ii) The “if” part is clear by the convexity of $x\mapsto 1/(1-\Lambda(x))$. To show the “only if” part, suppose (19) is convex in $x$. Let $\bar{x}=\inf\{x\in\mathbb{R}:\Lambda(x)=0\}.$ Right-continuity of $\Lambda$ yields that $\Lambda(\bar{x})=0$. We first prove $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in(-\infty,\bar{x})$. Suppose for contradiction that
@@
Next, we show that $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in\mathbb{R}$. Because $x\mapsto 1/(1-\Lambda(x))$ is decreasing, it suffices to show that $x\mapsto 1/(1-\Lambda(x))$ is continuous at $\bar{x}$ if $\bar{x}<\infty$. Suppose for contradiction that $\Lambda(\bar{x}-)>0$. Because of (22), there exists $a_{1}\in(-\infty,\mathrm{ess\mbox{-}sup}(X))$, such that $\mathbb{E}[a_{1}+(X-a_{1})_{+}/(1-\Lambda(\bar{x}-))]>\bar{x}.$ It is clear that $\mathbb{E}[(X-a_{1})_{+}]>0$ and thus
@@
$<\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]=\lim_{x\uparrow\bar{x}}\left\{\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda((x+\bar{x})/2)}(X-a_{1})_{+}\right]\vee\frac{x+\bar{x}}{2}\right\}.$
@@
$\mathbb{E}\left[\frac{a+b}{2}+\frac{1}{1-\Lambda\left(\frac{x+y}{2}\right)}\left(X-\frac{a+b}{2}\right)_{+}\right]\vee\frac{x+y}{2}$
$=\left\{\frac{a+t}{2}+\frac{1-\Lambda(x)}{1-\Lambda\left(\frac{x+y}{2}\right)}\left(t-\frac{a+t}{2}\right)\right\}\vee\frac{x+y}{2}>t.$
@@
For the RU formula of $\Lambda$-ES, Theorem 4 indicates that we do not guarantee joint convexity (or quasi-convexity) of the objective (19) in $(a,x)\in\mathbb{R}^{2}$ or $(a,x,X)\in\mathbb{R}\times\mathbb{R}\times L^{\infty}$ unless $\Lambda$ is a constant (i.e., $\mathrm{ES}_{\Lambda}$ is an ES). Nevertheless, the mapping (19) is convex in $x\in\mathbb{R}$ when the function $x\mapsto 1/(1-\Lambda(x))$ is convex. Two examples satisfying this condition are: $\Lambda$ is a constant (i.e. the usual ES) and $\Lambda(x)=(e^{ax}+1)^{-1}$ for a>0a>0.444The example $\Lambda(x)=(e^{ax}+1)^{-1}$, $a>0$, $x\in\mathbb{R}$, provides a suggestion of a non-trivial choice of the $\mathrm{ES}_{\Lambda}$ risk measure to use in practice. By Theorem 4, an obvious advantage of such a choice is that it makes the objective of a practical $\Lambda$-ES-based optimization problem convex in the variable $x$. Moreover, Theorem 4 shows that $\mathrm{ES}_{\Lambda}$ has a similar feature to a Bayes risk measure introduced by Embrechts et al. (2021), as it can be represented as the minimum of an expected loss function with an additional transformation. The corresponding minimizer is the interval of the left- and right-quantiles at the level of $\Lambda(x^{*})$ instead of $\mathrm{VaR}_{\Lambda},\mathrm{VaR}^{+}_{\Lambda}$].
@@ 5.2 Optimization with Lambda ES
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. We demonstrate general ideas of solving optimization problems with $\Lambda$-ES as a constraint or an objective. For $n\in\mathbb{N}$, let $\mathbf{L}=(L_{1},\dots,L_{n})\in(L^{\infty})^{n}$ represent a vector of losses, $\bm{\theta}\in\Theta$ represent a decision variable, where $\Theta$ is a convex set of actions, and $f:\Theta\times\mathbb{R}^{n}\to\mathbb{R}$ represent a loss function. For a typical example in finance, we can use $\mathbf{L}$ as the vector of losses from multiple assets and $\bm{\theta}$ as a portfolio weight vector.
First, we are interested in a problem where the decision maker minimizes an objective risk measure $\rho:L^{\infty}\to\overline{\mathbb{R}}$ of the aggregate loss $f(\bm{\theta},\mathbf{L})$, guaranteeing that the $\Lambda$-ES of the total loss does not exceed a pre-specified value $\ell\in\mathbb{R}$. Namely, we consider the following optimization problem:
@@
Proposition 4 implies that an optimization problem with a $\Lambda$-ES constraint of level $\ell\in\mathbb{R}$ can be equivalently converted to a problem with the constraint on $\mathrm{ES}_{\Lambda(\ell)}$ of the same level $\ell$. ES-constrained optimization problem has been studed extensively in the literature (see e.g., Krokhmal et al., 2002).
Another natural question to study is how to minimize $\Lambda$-ES as an objective:
@@
The result below provides general insights into solving $\Lambda$-ES-based optimization problems. In specific problems, we may also consider some constraints along with the problem (25), which does not change the nature of our equivalence result. Define the mapping $\widetilde{T}_{\Lambda}:\Theta\times\mathbb{R}\times\mathbb{R}\times(L^{\infty})^{n}\to\overline{\mathbb{R}}$ as
@@ 6.2 Dominance of ES and Lambda ES
One can check that $\rho$ is quasi-convex, law invariant, and $\rho\geqslant\mathrm{VaR}^{+}_{0}$. However, the condition $\rho\geqslant\mathbb{E}$ fails. Therefore, $\mathbb{E}$ is not the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}^{+}_{0}$ and thus (10) in Theorem 1 fails for $\alpha=0$.
@@
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law invariant}\}.$ (28)
@@
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law invariant}\}.$ (29)
@@
$\mathrm{ES}_{\alpha}=\min\{\rho:\underline{L}^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law invariant}\}.$ (30)
@@
The proofs of (28), (29), and the case of $\alpha\in(0,1)$ in (30) follow directly from those for Theorem 1 whose arguments still hold on $L^{1}$. We only need to prove (30) for $\alpha=0$.
@@
With Theorem 5, we obtain the following result, which extends the domain in Theorem 2 from $L^{\infty}$ to $L^{1}$ or $\underline{L}^{1}$.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1]$, the smallest quasi-convex and law-invariant risk measure on $L^{1}$ dominating $\mathrm{VaR}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR_Λ and $\rho$ is quasi-convex and law invariant}.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, the smallest quasi-convex and law-invariant risk measure on $L^{1}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law invariant}.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to[0,1)$, the smallest quasi-convex and law-invariant risk measure on $\underline{L}^{1}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law invariant}.
@@
6.3 Dual representation and RU formula for Lambda ES
In this section, we restate the dual representation (Theorem 3) and the RU formula (Theorem 4) for $\Lambda$-ES on $L^{1}$. The proofs of both results below follow from the same arguments as those for Theorems 3 and 4 by replacing $L^{\infty}$ with $L^{1}$.